An investor has a total of $26,400 deposited in 3 different accounts, which earn annual interset rates of 6%, 5%, 4%. The amount deposited in the 6% account is twice the amount in the 5% account. If the 3 acccounts earn a total annual nterest of $1331. How much money is deposited in each account?
Answer by stanbon(48535)

An investor has a total of $26,400 deposited in 3 different accounts, which earn annual interset rates of 6%, 5%, 4%. The amount deposited in the 6% account is twice the amount in the 5% account. If the 3 acccounts earn a total annual nterest of $1331. How much money is deposited in each account?
--------------
Quantity Eq: x + y + z = 26400
Quantity Eq: x - 2y + 0 = 0
Interest Eq: 0.06 + 0.05y + 0.04z = 1331
----------------------------------------
solve the system by any means you know to get:
----------
x = 11000 (amt. invested at 6%)
y = 5500 (amt. invested at 5%)
z = 9900 (amt. invested at 4%)
===================================
Chers,
stan H.